When Ricky was 10 years old, his father was 45.How many years must pass for Ricky to be 5 years older than 1/5 as old as his father?
Found 2 solutions by stanbon, chana:
Answer by stanbon(75887)

When Ricky was 10 years old, his father was 45.How many years must pass for Ricky to be 5 years older than 1/5 as old as his father?
-----
(10+x) = (1/5)(45+x)+5
---
5 + x = (1/5)(45+x)
25+5x = 45+x
4x = 20
x = 5
============
In 5 years dad will be 50
In 5 years son will be 15
----
Cheers,
Stan H.
============
Cheers,
Stan H.

Answer by chana(69)

When Ricky was 10 years old, his father was 45.How many years must pass for Ricky to be 5 years older than 1/5 as old as his father?
sol:
consider it as now,ricky = 10
father= 45
after x years ricky = 10 + x ------ (1)
in terms of father's age ricky = 1/5*(45+x) + 5 ------(2)
after x years ricky will be 5 years older than 1/5 of fathers
equate 1 & 2
you will get x = 5.